# Boosting Safe Human-Robot Collaboration Through Adaptive Collision Sensitivity

### v1

@@ V Experiments and Results
This Section describes the experiments done. We performed several real-world experiments using our setup that mimics possible collision scenarios in human-robot cooperation scenario. All data collected during our experiment are available at https://osf.io/9sqt3/.
V-A Experiments Settings
As stated in Section II-B, the robot performed a task inside our controlled environment with thee buckets ( each) that acted as obstacles (simulation of a collision with the human arm). For these obstacles, the robot was controlled to stop immediately after colliding with the robot skin at the moment a collision was detected. Consequently, the robot remains stopped for . If there is no longer a collision after the waiting period, the robot continues its task. If a new collision occurs at the same place (usually the bucket swings back to the robot), the robot stops again. After several collision detections (the number depends on skin sensitivity), the robot moves to the next location of its task. On top of the buckets, there is a measuring device located on the table, which simulates a clamping scenario between the robot and human operator. In case of a collision with this device, the robot is stopped again for , but now the robot moves back to the beginning of the current phase and then continues to the next waypoint of the task. The reason is that the measuring device cannot move out of the way of the robot, and thus the robot would stay in collision indefinitely.
The robot was controlled in joint velocity move so that the end-effector moves . However, we had to make the movement towards the quasi-static collision half the speed (), because otherwise the internal safety controller of the robot always acted sooner than our robotic skin-enabled controller. We tested 7 different threshold settings and repeated each one of them five times, resulting in 35 total runs of the task (with more than 65 collisions among all the runs).
V-B Task Performance
The first and, arguably, the most straightforward metric to evaluate the effect of skin sensitivity and the sensitivity setting method is the total time needed to complete the task. The results are shown in Fig. 5. We can see that, as expected, the more sensitive the skin, the longer the time required to complete the task. The difference between the most sensitive setting (Uniform 0) and the least sensitive one (Uniform 4) is, in average, more than , which is more than speed up.
The biggest reason for the difference in time between the uniform skin sensitivity settings is given by the number of collisions detected during the task. We provide a view on this parameter in Fig. 6. The first thing to notice is the total number of touches (detected collisions) for each sensitivity setting. The difference between Uniform 0 and Uniform 4 is, on average, more than 12 detections, which is a significant number. More specifically, the highest discrepancy in collision in different phases of the task is visible for collisions with buckets B1 and B3. For B3, the total number of collisions is low even in the most sensitive setting (Uniform 0), but is zero for Uniform 2, Uniform 3, and Uniform 4. For the case of bucket B1, the collision frequency starts at more than 8 collisions for Uniform 0, but drops to less than one for Uniform 4, showing the effect of skin sensitivity.
So far, we have talked about uniform (static) settings of skin sensitivity on robotic skin. However, a more interesting and promising method for setting skin thresholds is the dynamic change of thresholds for individual skin parts based on Eq. 2. We tested two methods. One that directly uses the force estimation equation from ISO/TS 15066 [6] and one that also utilizes the so-called Effective Mass (EF) of the robot (see Section IV). We can again use Fig. 5 and Fig. 6. In both of these, the two dynamic policies lie between Uniform 1 and Uniform 2, while the effective mass version is slightly more efficient (both in terms of shorter total time and lower number of collisions with the obstacles).
It may seem that adaptive policies are worse than some of the uniform ones. However, it is necessary to take into account the safety and forces exerted by the robot using different skin sensitivity. Unfortunately, our force measuring device is not suitable to measure transient collisions, and thus we can confidently assess safety only using quasi-static collisions. In Fig. 7 we show averaged (over 5 repetitions) force evolutions measured during movement in the negative z-axis of the robot. We can see that the more sensitive the threshold, the lower the mean maximal force and the smaller the SD between individual repetitions. The same is shown in Table I, together with the forces measured during experiments with dynamic strategies. We also provide estimates from both dynamic policies that are used to select the sensitivity threshold.
Force [N] 1 SD | |
Uniform 0 | 45.5 4.93 |
Uniform 1 | 59.6 20.4 |
Uniform 2 | 84.6 36.5 |
Uniform 3 | 87.6 44.6 |
Uniform 4 | 93.6 44.8 |
Dynamic HMM (measured) | 54.0 26.1 |
Dynamic EF (measured) | 61.8 27.8 |
Dynamic HMM (estimated) | 166.5 1.67 |
Dynamic EF (estimated) | 78.0 0.86 |
There are multiple things to reason about. First, we can see that the threshold sensitivity influences the exerted force (both in amplitude and variance). Then, the dynamic policies both overestimate the real forces, but the superiority of using Dynamic EF is easily notable. The policy that uses directly the equation for force from ISO/TS 15066 estimates the force as , which is actually even over the official limit for quasi-static contacts (). Thus, it sets the most sensitive threshold (see also Table II) which results in low real force (). On the other hand, the policy that uses EF of the robot estimates the force as (less than half compared to the first dynamic method). That results in using the third most sensitive threshold (see again Table II) with real force measured of , which is considerably closer to the estimate. We can also look in Table II where the thresholds set by the dynamic methods for all collision places are shown. We can see that Dynamic HMM method sets the most sensitive thresholds everywhere, except for bucket B3. In contrast, the Dynamic EF sets the most sensitive threshold only for the first bucket and uses different thresholds for other obstacles.
Impact Position [-] | ||||
B1 | B2 | QS | B3 | |
Dynamic HMM | 0 | 0 | 0 | 2 |
Dynamic EF | 0 | 1 | 2 | 3 |
VI Conclusion, Discussion, and Future Work
We introduced different strategies to set skin sensitivity thresholds—five uniform and two dynamic. We tested the policies in the real world with our custom environment that enabled us to simulate both transient and quasi-static collisions in a HRI task. We measured the impact forces in a quasi-static (clamping) scenario using a certified measuring device. We showed that skin sensitivity has an influence on task execution time, the number of collision detections, and the forces exerted. We further demonstrate that using a dynamic setting while computing Effective Mass (EF) of the robot can estimate the exerted force more precisely than using a basic formula from the corresponding ISO standard [6], providing better task performance while preserving human safety.
@@ VI Conclusion, Discussion, and Future Work
In this work, only the simplest reaction to a collision that exceeds a pressure threshold was applied: stop the robot. Post-collision reactions can have several forms and may also depend on what information is available about the collision (see collision identification in [8]) and on the robot capabilities (e.g., whether torque control is available [7]). In the setup used here—with coarse isolation of the collision location in space, a pressure value reading, and a velocity control mode—one possibility is to implement an avoidance reflex that will move the colliding robot link in the opposite direction (see [15]). Artificial skins with higher spatial resolution make it possible to move away the collision point along the normal to the robot skin at that location [17].

### v2

@@ III Experiments and Results
We performed several simulated and real-world experiments using our setup that mimics possible collision scenarios in human-robot collaboration. All data collected during the experiments are available at https://rustlluk.github.io/adaptive-sensitivity/.
III-A Experiments Settings
As described in Section II-A, the robot performed a mock pick-and-place task in a controlled environment with three buckets ( each) suspended from the ceiling that acted as obstacles (simulation of a collision with the human arm). Collisions were detected and isolated and depending on the collision reaction strategy, the robot was commanded to stop or to continue task operation. If stop was triggered, the robot remained stopped for . If there was no longer a collision after the waiting period, the robot continued its task. If a new collision occurred at the same place (usually the bucket swinged back to the robot), the robot evaluted it, possibly stopping again. In addition to the buckets, there was an impact measuring device attached to the table, which simulated a clamping scenario between the robot and a human operator. As this contact was not transient, if a stop was triggered, the robot stopped for and then was commanded to continue the task.
The robots were controlled in joint velocity so that the end-effector moved with a predefined Cartesian velocity. However, we had to make the movement towards the quasi-static collision at max in order not to trigger the internal safety controller of the robot.
In simulation, we tested both robots with Cartesian velocity of 200, 400, and . For each velocity and each collision strategy (FACTORY, FIXED MASS, ADAPTIVE MASS) we ran 100 repetitions of the experiments. In the real world, we tested the UR10e robot with velocities of and and performed 10 repetitions for each velocity and collision reaction strategy.
III-B UR10e with AIRSKIN
We will start with the UR10e robot, for which we performed experiments in both the simulation (left side in Fig. 5) and the real world (right side in Fig. 5). The first, and arguably most important, results in both worlds is the difference in task completion time between different collision handling strategies for the Cartesian velocity of . As expected, the FACTORY mode, where every touch results in a stop of the task execution, is the slowest. It is followed by the FIXED MASS and ADAPTIVE MASS modes. In the real world, the difference between the FACTORY and ADAPTIVE MASS modes is, on average, more than 6 s (45% faster; 14.5 s vs. 8 s) while FIXED MASS is closer to ADAPTIVE MASS with 10 s. In the simulation, the times are closer to each other with 11.5, 9, and 7.5 s, respectively. The discrepancy is probably cased by the physics of the buckets in the simulation. We tried to set the parameters as closely to the real world as possible (see the video at https://youtu.be/UyOa--6Jpj4), but the behavior is still not identical.
The completion time of the task is correlated with the number of stops of the robot during the task, shown in the bottom part of the Fig. 5. We can see that the trend in differences between the individual modes is the same as for the task time. In the real world, the robot stopped using ADAPTIVE MASS only half of the times (on average) compared to the FACTORY mode (4 vs. 8 stops), while FIXED MASS is again between them with 6 stops. This behavior is supported by Fig. 1, where, mainly for bucket B2, the ADAPTIVE MASS strategy is more permissive. The average number of stops is lower in the simulation with 4.5, 3.5, and 2.5 for FACTORY, FIXED MASS and ADAPTIVE MASS, respectively.
With lower velocities, there is basically no difference between ADAPTIVE MASS and FIXED MASS, as the velocities are too slow to exceed the stop thresholds for either of the strategies. We can see that for the velocity of the robot stopped on average only once during the task for both modes and both worlds. And for the velocity of it actually never needed to stop. In the FACTORY mode, the task time (and number of stops) is high. For example, in the real world with the velocity of the difference in task time is 21 s vs 6.3 s.
The results show that using ADAPTIVE MASS or FIXED MASS makes task completion faster with fewer interruptions. There is thus a clear increase in productivity. We need to verify that this does not compromise safety. In other words, we need to verify that the impact force estimates are conservative. Therefore, we utilized the CBSF-75-Basic certified measuring device and measured the forces excerted by the robot during quasi-static contact (the limit based on ISO/TS 15066 is 140 N). The results are shown in Fig. 6. The histogram is created from all 60 measurements in the real world and all were measured with a velocity of . We can see that more than 90% of the forces is lower than the force estimated using ADAPTIVE MASS (red line; 175 N). The measuring device is sensitive to the “accuracy” of the impact (it should be in the middle and in the direction of the spring of the device) and thus the variance of the measurements is high (from 40 N to 215 N), but the majority of forces are under 160 N. In a real deployment scenario, such empirical measurements would be used to adapt the dynamic thresholds, making sure the prescribed limits are always respected.
As described in Section II-D we also attempted to measure the impact force exerted during transient contacts, using a pendulum approximation. We extracted the maximum angles after collision from the real videos of a protractor behind the bucket B1. We used only values from experiments with FACTORY strategy, as it always stopped and it was easier to extract the angles. The measured angles were and for the velocities of and , respectively. The forces estimated using ADAPTIVE MASS near bucket B1 for the velocities are 190 and 280 N, respectively. Pendulum forces computed using Eq. 8 for different contact durations are shown in Table I. It is complicated to estimate the correct , but from the recordings we know that it is more than one frame and less than two, that is, 40-80 ms. We can see in Table I that the forces for 40 and 50 ms are low. However, we must take into account that these are average forces exerted throughout the duration of contact. There will probably be a higher initial force peak exerted for a shorter time duration. As both the bucket and AIRSKIN are slightly deformable, a contact duration of around 20-30 ms is a reasonable estimate. As we can see, the forces are still lower than the forces calculated using ADAPTIVE MASS.
| Contact Duration [ms] | ||||
| Velocity | 20 | 30 | 40 | 50 |
| 400 | ||||
| 600 | ||||
Based on the fact that both quasi-static and transient estimates are lower than the actual forces, we can say that ADAPTIVE MASS is effective and also still conservative, and thus safe.
III-C KUKA LBR iiwa 7
To demonstrate the general applicability of our method—it can work on any robot capable on collision detection and isolation, we also evaluated the task on a simulated KUKA LBR iiwa 7 robot. This robots does not have sensitive skin, but each of its joints is equipped with joint torque sensors that allow to detect and isolate collisions. We ran an adapted version of the same task with buckets at different places as the robot is smaller than the UR10e.
The results are shown in Fig. 7. We can see that the trends are similar to those for the UR10e. At the highest velocity () the task time is the highest for FACTORY mode (11 s) followed by FIXED MASS (10 s) and ADAPTIVE MASS (8 s), resulting in a speedup of about 30% when using ADAPTIVE MASS. The number of stops again corresponds to the task time with 4.5, 4, and 3 touches, respectively. Interestingly, at lower velocities the ADAPTIVE MASS is slower (and stops more) than FIXED MASS. It shows that FIXED MASS can actually underestimate the situation in some cases and the force computed using ADAPTIVE MASS is higher, which may be an important aspect of safety.
IV Conclusion, Discussion, and Future Work
We introduced three different collision handling strategies: One static and commonly used in the industry, where every detected collision triggers a stop of the robot, and two dynamic strategies that are calculating estimated contact force on the fly to choose how to react to the collisions—stop or continue operation. We tested the strategies using two robots in simulation (UR10e with artificial skin AIRSKIN and KUKA LBR iiwa 7 with joint-torque sensors) and with UR10e robot in the real world. The results show that when utilizing one of our dynamic methods (ADAPTIVE MASS), which calculates the estimated force using a standardized equation from ISO/TS 15066 [6] and Effective Mass (EF) of the robot, the task time can be significantly lower (up to 70% with the velocity of or 45% with , for the UR10e robot). The strategies also result in different number of interruptions during the task, which can be crucial for the health of the robot and its components.
The results are consistent for the two different robots we used, each of them using a different collision detection and isolation strategy (sensitive skin vs. joint torque sensing). Other collision detection methods are used by collaborative robots (a force/torque sensor in the robot base or using motor currents). Our method is potentially applicable to any of them, provided that collision detection and isolation (which robot link collided) is available. For the strategy that increases productivity the most (ADAPTIVE MASS), a code must be developed to estimate the effective mass of every robot link online.
We compared the collision forces estimated by our method with empirical measurements to make sure that the increase in productivity does not compromise safety. Some empirical measurements will always be indispensable before real deployment (see [18, 12] for extensive empirical measurements on the two robots used in this work).
@@ IV Conclusion, Discussion, and Future Work
In this work, only the simplest reaction to a collision that exceeds a pressure threshold was applied: stop the robot. Post-collision reactions can have several forms (see, for example, [19]) and may also depend on what information is available about the collision (see collision identification in [7]) and on the robot capabilities (e.g., whether torque control is available [8]). In the setup used here—with coarse isolation of the collision location in space, a pressure value reading, and a velocity control mode—one possibility is to implement an avoidance reflex that will move the colliding robot link in the opposite direction (see [14]). Artificial skins with higher spatial resolution make it possible to move away the collision point along the normal to the robot skin at that location [20].
Our method is usable for contacts with parts of the robot that are not sharp or dangerous otherwise. However, the forces and possible injuries change based on the colliding part of the robot (see [21]), which should be incorporated into the pipeline.